### v2

# Enhancing Multi-Agent Collaboration with Attention-Based Actor-Critic Policies

Abstract
This paper introduces Team-Attention-Actor-Critic (TAAC), a reinforcement learning algorithm designed to enhance multi-agent collaboration in cooperative environments. TAAC employs a Centralized Training/Centralized Execution scheme incorporating multi-headed attention mechanisms in both the actor and critic. This design facilitates dynamic, inter-agent communication, allowing agents to explicitly query teammates, thereby efficiently managing the exponential growth of joint-action spaces while ensuring a high degree of collaboration. We further introduce a penalized loss function which promotes diverse yet complementary roles among agents. We evaluate TAAC in a simulated soccer environment against benchmark algorithms representing other multi-agent paradigms, including Proximal Policy Optimization and Multi-Agent Actor-Attention-Critic. We find that TAAC exhibits superior performance and enhanced collaborative behaviors across a variety of metrics (win rates, goal differentials, Elo ratings, inter-agent connectivity, balanced spatial distributions, and frequent tactical interactions such as ball possession swaps).
Index Terms:
Reinforcement Learning, Multi-Agent Systems, Actor-Critic, Attention Mechanism, Centralized Training and Execution (CTE)
I Introduction
When human organizations address a large cooperative problem or a task, they scale the number of individuals at the expense of having to organize and coordinate them. Over thousands of generations, we have honed our ability to organize many individuals to collaborate at scale to reach a common goal, and woven it into our largest architectural and organizational achievements. As we integrate reinforcement learning into real-world applications, we need multi-agent systems that efficiently scale collaborative action to large numbers of agents to address broader needs and problems. In this paper, we study methods to manage cooperative multi-agent environments or systems. Our goal is to teach groups of AI agents strong collaborative skills that allow them to act as a team in a scalable fashion, thereby tackling complex problems and challenges efficiently.
The best methods for training a team to collaborate in a cooperative environment remains an open question. The main paradigms that have been used to learn collaboration across agents in cooperative MARL are Decentralized Training/Decentralized Execution (DTDE), Centralized Training/Decentralized Execution (CTDE), and Centralized Training/Centralized Execution (CTCE). In DTDE agents operate independently without shared information both during training and inference while in CTDE information is shared during training (usually to estimate values) but not when agents select actions or during inference. Finally, CTCE shares information during execution, which allows delegation of control to a single controller which determines the joint-action of all agents. DTDE and CTDE schemes present challenges due to the need to decide whether and how to share information across agents; these modeling choices are critical for team collaboration. [^1] Decentralized execution schemes can also suffer from convergence problems due to the inherent non-stationarity that arises when multiple agents learn and adapt concurrently. The CTCE single controller avoids these difficulties but poses another problem: the full joint-action space under a CTCE scheme is the product set of individual agent action sets and so its size grows exponentially in the number of agents. Effective team collaboration would seem to hinge on a robust search of the joint-action space for complimentary actions across agents, and so a CTCE scheme imposes scalability constraints on collaborative activity.
To address these challenges, we introduce and test an RL algorithm that leverages shared information to foster inter-agent collaboration. Our proposed model, Team-Attention-Actor-Critic (TAAC), is a CTCE scheme that incorporates an attention mechanism directly into the actor to model an information-sharing “conversation” between multiple agents which enables them to dynamically query teammates during decision-making. This allows our agents to manage the joint-action space in a scalable fashion. We further introduce a loss that encourages agents to learn diverse yet complementary roles, enhancing collaboration. We evaluate TAAC in a simulated soccer environment, an ideal testbed for teamwork and spatial coordination, and benchmark its performance against PPO (as an example of DTDE) and MAAC (as an example of CTDE) using both standard metrics (e.g., win rate, Elo) and multiple indicators of collaboration (e.g., pairwise distance, possession swaps, connectivity).
II Related Work
In cooperative MARL environments, CTCE seems a natural choice to maximize collaborative potential, especially given the robustness of modern information sharing capabilities which would seem to allow for relatively costless message sending where agents either do not wish to or cannot restrict information sharing. A number of recent papers on MARL have proposed managing agent collaboration by leveraging modern deep learning architectures. Within the CTCE paradigm, [1] and [2] break down the full joint-action space by requiring agents to select actions sequentially and treating MARL as a sequence modeling problem by applying transformers. [3] and [4] apply attention to deal with high-dimensional action spaces in a value-based context (to variants of deep $Q$-learning). By contrast, we focus on simultaneous action selection in a policy-based algorithmic context. 222[1] and [2] would converge to a Markov Perfect Equilibrium for a stochastic game with within-period sequential actions by the agents; there are no guarantees that the equilibrium set for such a game would map well onto the equilibrium set for a game with true simultaneous choices.
Beyond this, some CTDE algorithms have applied attention to model communication between agents, particularly [5] and [6] which use attention mechanisms to learn when and with whom to communicate. The former models communication groups dynamically via attention gating. The latter models targeted communication using attention mechanisms, enabling message exchange between specific agents. Finally and perhaps most relevant for our work, MAAC [7] learns a centralized critic with a soft attention mechanism that dynamically selects agents to attend to at each time step during training.  [^3] Other CTDE schemes have explored different approaches to deal with information within multi-agent environments by modeling communication and information sharing between agents via a combination of heuristic rules (e.g. [8], [9]), and explicitly modeling message selection (e.g. [10]), message timing (e.g. [11] extending [12]), or message recipients (e.g. [13], [14]).
These CTDE methods are particularly appropriate for teaching agents to manage environments collaboratively in which information sending is problematic, e.g. because agents are physically separated, because messages are costly to send, or because there are incentives to maintain private information. At the same time, if information sending is not the main constraint  [^4] Examples of such environments might be sports teams, simulated environments, digital games, or any environment in which a slight time lag to build a full picture of observations across agents is acceptable. then the focus on decentralized execution leads to more complicated methods than might be needed  [^5] Predefined or hueristic communication architectures restrict communication and hence restrain potential collaboration among agents. A potential critique of [8] and [9] is that use of an aggregative operation across agents has the potential to eliminate differences among neighbouring agents and result in loss of important information. , likely underestimates the extent to which collaborative action is possible, and relegates it to no more than an implicit, emergent property during inference.
In this paper, we build upon [7] and [15]. We extend their work by incorporating an attention mechanism in the actor and a novel loss which allows agents to receive information in a selective manner. This allows us to let agents collaborate explicitly during inference. It functions by allowing an agent to pose an open-ended question to its peers. For example, in a soccer game scenario, a player with the ball may query, ‘Who should I pass to?’ Agents which are not nearby opponents may be identified as optimal recipients, and their responses inform the passing decision. We additionally develop a novel loss (named conformity loss) which encourages diverse actions in agents, enhancing collaboration in certain environments. We have named this model Team-Attention-Actor-Critic (TAAC).
To test our algorithm, we train it in a simulation of a soccer environment. In this setting, agents control individual players on a team competing in a virtual soccer match. We present our results by comparing the performance of our approach against state-of-the-art algorithms using a combination of traditional performance metrics (such as win rates and ELO ratings) and collaboration metrics (such as inter-agent connectivity and pairwise distance).
This paper is organized as follows. In Section 3, we introduce the notation and provide the necessary background. Section 4 details our proposed algorithm along with its implementation. In Section 5, we describe the experimental setup and the evaluation metrics employed. Section 6 presents the experimental results, while Section 7 offers an in-depth analysis and discussion of these findings. Finally, Section 8 concludes the paper and outlines directions for future work.
III Background and preliminaries
Our strategy to model collaboration will build upon actor-critic algorithms. Below, we introduce this algorithm along with the principles and notation needed to formalize it. Following this section, we provide a complete description of our architecture.
In reinforcement learning, control problems with multiple agents can be modelled as Markov Decision Games (MDG), which are defined by:
  • A set of $n$ agents;
  • A set of states $S$ that describe the current common environmental conditions facing the agents;
  • A collection of actions $A(s)=\prod_{i}A_{i}(s)$ for $s\in S$ where $A_{i}(s)$ are the actions agent $i$ can take in state $s$ and:
    $\vec{a}=\langle a_{1},a_{2},\ldots,a_{n}\rangle\in A(s);$
  • Transition probabilities $p(s^{\prime}\mid\vec{a},s)$ for transitioning from state $s$ to state $s^{\prime}$;
  • A function:
    $\vec{r}(s^{\prime},\vec{a},s)=\langle r_{i}(s^{\prime},\vec{a},s)\rangle_{i=1,\ldots,n}$
    that supplies the immediate rewards for each agent associated with this transition.
[…]
$\vec{z}_{\backslash i}=\langle z_{1},\ldots,z_{i-1},z_{i+1},\ldots,z_{n}\rangle$
and we will write $\vec{z}_{i}$ to mean the pair $(z_{i},\vec{z}_{\backslash i})$.
In environments that take place over a finite number of discrete periods $T$, the sequence of periods in which the agents participate is referred to as an episode. At each time step, each agent makes an observation, $o_{i,t}=O_{i}(s_{t})$, of the global state, which encapsulates all available information for agent $i$; the vector of such observations across agents is $\vec{o}_{t}=\langle o_{i,t}\rangle_{i=1,\ldots,n}$ The goal of each agent is to learn a policy which describes the probability that a trained agent should take action $a_{i}$ given an observation $o_{i,t}$, to maximize its individual sequence of rewards throughout an episode: $\sum_{t=0}^{T-1}\gamma^{t}r_{i}(s_{t+1},\vec{a}_{t},s_{t})$. Here $\vec{a}_{t}$ and $s_{t}$ are the agents’ actions and the common state at time $t$ and $\gamma\in[0,1]$ is the discount factor on future rewards. In general, policies could differ across agents; we will focus on the case where the policy is shared. Any differences in decision-making across agents are modelled using their observations of the state. Given this, we write these policies as $\pi(a_{i}|o_{i,t})$.
III-A Policy Gradients
Policy gradient methods model this policy directly, usually via a function approximator with parameters $\mathbf{u}$; we indicate the dependence of the policy on these parameters by writing $\pi_{\mathbf{u}}(\cdot)$. [^6] By contrast, value-based methods (e.g. $Q$-learning, [16]) estimate the state-conditional value of actions and then use those values to build a policy. To learn $\mathbf{u}$ for $\pi_{\mathbf{u}}(\cdot)$, these methods use an estimate of the policy gradient ([17], [18], [19]) to update the parameters in the direction of increase in the expected sum of discounted rewards. In the case of $n$ agents with shared policies, taking the Monte Carlo approach of using the full return, the policy gradient may be written as:
$\nabla_{\mathbf{u}}J(\mathbf{u})=\mathbb{E}_{\tau}\Biggl{[}\sum_{i=1}^{n}\sum_{t=0}^{T-1}\nabla_{\mathbf{u}}\ln\pi_{\mathbf{u}}\bigl{(}a_{i,t}\mid o_{i,t}\bigr{)}\,[G_{i,t}(\tau)-b(o_{i,t})]\Biggr{]},$ (1)
where $\tau$ is a trajectory of states and actions whose distribution is dependent on $\pi_{\mathbf{u}}(\cdot)$ and on unknown environmental dynamics:
$\tau=s_{0},\vec{a}_{0},s_{1},\vec{a_{1}},\ldots,s_{T-1},\vec{a}_{T-1},s_{T}$
and $G_{i,t}(\tau)$ is the sum of discounted rewards derived from trajectory $\tau$ from a given timestep $t$ onwards (the causal return) for a given agent $i$:
$G_{i,t}(\tau)=\sum_{k=t}^{T-1}\gamma^{\,k-t}\,r_{i}\bigl{(}s_{k+1},\,a_{k},\,s_{k}\bigr{)}.$ (2)
Finally, $b(\cdot)$ is a baseline subtracted from the return to avoid high variance in the estimates of $\nabla_{\mathbf{u}}J(\mathbf{u})$ across episodes driven by use of $G_{i,t}(\tau)$. [^7] An arbitrary function of the state can be introduced into the policy gradient equation without changing the expectation – see [17] for details. If $b(\cdot)$ is allowed to depend on agent actions then this will bias the estimate of the gradient.
Given sampled episodes, this estimate of the policy gradient is used to take steps in the direction of $\nabla_{\mathbf{u}}J(\mathbf{u})$ via gradient ascent to reach an optimal policy.
III-B Actor-Critic Algorithms
A common choice for the baseline function $b(\cdot)$ is an estimate of the long-term value, for example, the state-value function, $V(o_{i,t})$, which predicts the return to an agent given an observation $o_{i,t}$ (see [20]). If we model the state value function as a differentiable function of parameters $\mathbf{w}$ and write $V_{\mathbf{w}}(\cdot)$ then this yields the following baseline:
$b(o_{i,t})$ $=V_{\mathbf{w}}(o_{i,t})$
$=\sum_{a\in A_{i}}\pi_{\mathbf{u}}(a|o_{i,t})Q_{\mathbf{w}}(o_{i,t},a)$
where the second equality follows from the usual definitions of the state-value function $V_{\mathbf{w}}(\cdot)$ and the state-action value function $Q_{\mathbf{w}}(\cdot)$. We will use this form of the baseline below to model multi-agent structure.
The policy is referred to as the actor and the long-term value estimate as the critic, and the resulting model is called an actor-critic algorithm. The actor’s parameters are learned through the policy gradient estimates and the critic’s parameters are learned through value-based methods (e.g. gradient descent on the difference between the critic’s predicted state value and a target value, which can be either the observed $G_{i,t}(\tau)$ or a bootstrapped value, see [21]).
IV Team-Attention-Actor-Critic (TAAC)
IV-A Attention Mechanism
We extend [7] by introducing a separate attention mechanism in the actor. The intuition behind doing this is to allow agents to use the internal representations of other agents directly when making choices, with the idea being that enabling this is analogous to allowing agents to “put themselves into the shoes” of their colleagues and thereby giving them insight into what other agents will choose. If they build an understanding of what their colleagues will choose then they may be able to create collaboration by selecting actions that dovetail well.
In general, attention works as follows (see [22]). An initial input $z$ is suitably embedded as a sequence of vectors $M(z)=\{m_{1}(z),\ldots,m_{n}(z)\}$ and multiplied with learned matrices $W_{Q}$, $W_{K}$, and $W_{V}$. Attention is then computed as:
$A(M(z)|W)=\text{Softmax}\left(\frac{M(z)W_{Q}(M(z)W_{K})^{T}}{\sqrt{d_{K}}}\right)M(z)W_{V}$ (3)
where $W=(W_{Q},W_{K},W_{V})$ and $d_{K}$ is the dimensionality of the $M(z)W_{K}$ matrix. [^8] The subscripts refer to the usual presentation of $W_{Q}$, $W_{K}$, and $W_{V}$ as the query, key, and value matrices, respectively. Note that in all other sections of the paper $Q(\cdot)$ will refer to the state-action value function. Division by $d_{K}$ is done to scale the values for numerical stability.
In a natural language application, the input might be a sequence of embedded tokens (e.g. words). In our application, these embedded input tokens will represent individual agent observation and/or action information. We apply multiple attention heads in parallel, each with its own set of weight matrices. The updated outputted sequence of tokens from these heads are concatenated by agent to create an updated set of agent-specific embedding information. Specifically, given $h$ attention heads, for agent $i$ we gather updated agent-specific embedding information by concatenating:
$E_{i}(z|W)=\{A_{i,j}(M(z)|W_{j})\}_{j=1}^{h}$ (4)
where $A_{i,j}(\cdot|W_{j})$ refers to the values in the $i$th row of the $j$th head’s output.
These tokens are individually forwarded through the rest of the architecture to generate either a distribution over actions (in the actor) or a state-action value estimate (in the critic); we describe both below.
Figure 1: Architecture of Actor. Note that the output is a set of $n$ distributions over actions, one for each agent.
IV-B Actor Attention Architecture
We model the policy with multiple output layers that share parameters but take different inputs, one for each agent. To incorporate attention into the policy, we condition it on the observations of all agents and write it as $\pi_{\mathbf{u}}(\vec{o})\in\prod_{i}\Delta(A_{i})$ so that $\pi_{\mathbf{u}}(\vec{o})=\langle\pi^{(1)}_{\mathbf{u}}(\vec{o}),\ldots,\pi^{(n)}_{\mathbf{u}}(\vec{o})\rangle$. Thus, $\pi_{\mathbf{u}}(\cdot)$ generates $n$ distributions, one over each action set $A_{i}$, and $\pi^{(i)}_{\mathbf{u}}(a_{i}|\vec{o})$ returns the probability that agent $i$ takes action $a_{i}\in A_{i}$. [^9] We focus on the case where $A$ is discrete here. The forward pass of the parameterized policy network is as follows.
  1. 1.
    For $i=1,\ldots,n$ the observation $o_{i}$ is passed through an appropriate embedding layer, say $m^{\pi}_{1}(\cdot|u_{m^{\pi}_{1}})$, where $u_{m^{\pi}_{1}}$ are associated parameters, to generate initial embeddings which are formed into the matrix:
    $M_{\pi}(\vec{o}|u_{m^{\pi}_{1}})=\left(\begin{array}[]{c}m^{\pi}_{1}(o_{1}|u_{m^{\pi}_{1}})\\ \vdots\\ m^{\pi}_{1}(o_{n}|u_{m^{\pi}_{1}})\end{array}\right).$
  2. 2.
    These embeddings are passed into a layer of $h_{\pi}$ attention heads which yields:
    $A_{j}(M_{\pi}(\vec{o}|u_{m^{\pi}_{1}})|u_{A_{j}}).$
    where $u_{A_{j}}=(u_{A_{j},K},u_{A_{j},Q},u_{A_{j},V})$ are the parameters for the $j$th actor attention head.
  3. 3.
    For every agent $i=1,\ldots,n$ we gather updated agent-specific embedding information by concatenating as described in equation 4, which we write in terms of the input observation vector as $E_{i}(\vec{o}|u_{m_{1}^{\pi}},u_{A})$ where $u_{A}=(u_{A_{1}},\ldots,u_{A_{h_{\pi}}})$.
  4. 4.
    Each updated agent-specific embedding is passed through a set MLP layers, say $m^{\pi}_{2}(\cdot|u_{m^{\pi}_{2}})$ where $u_{m_{2}^{\pi}}$ are associated parameters, and a final softmax operation is applied to generate $n$ distributions over actions, one for each agent.
IV-C Critic Attention Architecture
In the critic, we incorporate both the observations and actions of other agents. Consequently, the critic’s state-action values are written as $Q_{\mathbf{w}}(\vec{o},\vec{a})\in\mathbb{R}^{n}$ so that $Q_{\mathbf{w}}(\vec{o},\vec{a})=\langle Q^{(1)}_{\mathbf{w}}(\vec{o},\vec{a}),\ldots,Q^{(n)}_{\mathbf{w}}(\vec{o},\vec{a})\rangle$ and $Q^{(i)}_{\mathbf{w}}(\vec{o},\vec{a})$ provides that state-action value estimate for agent $i$ taking action $a_{i}\in A_{i}$. The forward pass of the parameterized critic network is as follows:
  1. 1.
    For $i=1,\ldots,n$ the observation $o_{i}$ and action $a_{i}$ are passed through an appropriate embedding layer, say $m_{1}^{Q}(\cdot|w_{m_{1}^{Q}})$ where $w_{m_{1}^{Q}}$ are associated parameters, to generate initial embeddings which are formed into the matrix:
    $M_{Q}(\vec{o},\vec{a}|w_{m_{1}^{Q}})=\left(\begin{array}[]{c}m^{c}_{1}(o_{1},a_{1}|w_{m_{1}^{Q}})\\ \vdots\\ m^{c}_{1}(o_{n},a_{n}|w_{m_{1}^{Q}})\end{array}\right).$
  2. 2.
    These embeddings are passed into a layer of $h_{Q}$ attention heads which yields:101010in [7], their attention mechanism differs only by having the $V$ matrix be fixed for each attention head.
    $A_{j}(M_{c}(\vec{o},\vec{a}|w_{m_{1}^{Q}})|w_{A_{j}})$
    where $w_{A_{j}}=(w_{A_{j},K},w_{A_{j},Q},w_{A_{j},V})$ are the parameters for the $j$th critic attention head.
  3. 3.
    For every agent $i=1,\ldots,n$ we gather updated agent-specific embedding information by concatenating as described in equation 4, which we write in terms of the input observation and action as $E_{i}(\vec{o},\vec{a}|w_{m_{1}^{Q}},w_{A})$ where $w_{A}=(w_{A_{1}},\ldots,w_{A_{h_{Q}}})$.
  4. 4.
    Each updated agent-specific embedding is combined with the original embedding and passed into a set of MLP layers, say $m^{Q}_{2}(\cdot|w_{m_{2}^{Q}})$ where $w_{m_{2}^{Q}}$ are associated parameters, to generate $n$ state-action value estimates, one for each agent.
With all these steps in mind we may write the output of the policy for the $i$th agent as:
$\pi^{(i)}_{\mathbf{u}}(\cdot|\vec{o})=\mbox{Softmax}\left(m^{a}_{2}\left(\left.E_{i}(\vec{o}|u_{m_{1}^{\pi}},u_{A})\right|u_{m_{2}^{\pi}}\right)\right)$
where $\mathbf{u}=(u_{m_{1}^{\pi}},u_{A},u_{m_{2}^{\pi}})$ and the output of the critic for the $i$th agent:
$Q^{(i)}_{\mathbf{w}}$ $(\vec{o},\vec{a})$
$=m^{Q}_{2}\left(\left.m_{1}^{Q}\left(o_{i},a_{i}|w_{m_{1}^{Q}}\right),E_{i}\left(\vec{o},\vec{a}_{i}|w_{m_{1}^{Q}},w_{A}\right)\right|w_{m_{2}^{Q}}\right)$
where $\mathbf{w}=(w_{m_{1}^{q}},w_{A},w_{m_{2}^{q}})$, respectively.
We conclude the description of the architecture with two notes. First, we deliberately omitted the original embedding as a direct input to the post-attention layers of the actor. This design choice forces the actor to rely solely on the information provided by the attention blocks rather than bypassing them, ensuring collaborative policies. By contrast, in the critic architecture, we include the original embedding to provide additional information, as the critic does not require the same enforced reliance on collaborative signals as the actor does. Second, each of the $n$ outputs in both the actor and critic make use of all model parameters, meaning that there is no obvious way to factor either model into agent-specific policies and critics that do not share parameters and information. The actor and critics architectures are depicted in figure  1 and  2, respectively.
Figure 2: Architecture of Critic. Note that the output is a set of $n$ state-action values, one for each agent.
IV-D Multi-Agent Baseline
We adopt the approach taken in [15] and set the baseline in the policy gradient by adapting the traditional critic to the multiagent setting in the following way:
$b_{i}(\vec{o}|\vec{a}_{\backslash i})$ $=V(\vec{o}|\vec{a}_{\backslash i})$
$=\sum_{a_{i}\in A_{i}}\pi^{(i)}_{\mathbf{u}}\left(a_{i}\mid\vec{o}\right)\,Q^{(i)}_{\mathbf{w}}(\vec{o},\vec{a}).$
Here, baseline models the multiagent setting in the sense that it estimates the expected value for a specific agent, conditioning on the observations and actions of all other agents.
IV-E Conformity Loss
We further model and encourage collaboration by adjusting the actor loss function to promote diversity of actions among agents, under the assumption that effective collaboration arises when agents work on separate and diverse tasks. [^11] The value of diversity may not hold in all environments, but we found it to be a useful addition for our experiments. We call this adjustment conformity loss. It is computed by measuring the cosine similarity between each unique pair of updated agent-specific embeddings derived from attention layer outputs.
$\mathcal{L}$ $(u_{m_{1}^{\pi}},u_{A})$
$=\theta_{S}\max\left\{\frac{2}{n(n-1)}\sum_{i}\sum_{i^{\prime}\neq i}S_{ii^{\prime}}(u_{m_{1}^{\pi}},u_{A}),\theta_{B}\right\}$
where:
$S_{ii^{\prime}}(u_{m_{1}^{\pi}},u_{A})=\frac{E_{i}(\vec{o}|u_{m_{1}^{\pi}},u_{A})^{\top}E_{i^{\prime}}(\vec{o}|u_{m_{1}^{\pi}},u_{A})}{\|E_{i}(\vec{o}|u_{m_{1}^{\pi}},u_{A})\|\,\|E_{i^{\prime}}(\vec{o}|u_{m_{1}^{\pi}},u_{A})\|}$
and $\theta_{S}$ and $\theta_{B}$ are hyperparameters used to scale and bound the loss. This contribution to overall loss attains high values when agents’ attention layer outputs are very similar (indicating low diversity) and low values when they are not. [^12] In our implementation, we lower bound this loss at a minimum of $\theta_{B}$ to ensure that parameter gradients are directed toward an intermediate level of action diversity during learning.
V Experimentation
In this paper, we compare multiple algorithms against our novel architecture (TAAC) in a simulated soccer environment. In this environment, teams of 3 agents work together to score goals against an opposing team. A goal is scored when the ball is kicked into the opponent’s goal. Each team is composed entirely of agents running the same algorithm to ensure a fair comparison.
V-A Environment Structure
The simulation represents players as solid circles on a 2D plane capable of moving toward the ball and colliding with it. When a player collides with the ball while in mode “kicking”, the ball gains additional velocity. Additionally, the ball collides nearly elastically with the walls, preventing it from leaving the play area. Figure 3 illustrates the simulation environment.
Figure 3: Simulation Environment for Soccer
At each time step, every agent derives an individual, local observation from the global state. In our environment, the state consists of:
  • The relative position vector to each teammate.
  • The relative position vector to each agent on the opposing team.
  • The relative position vector to the ball.
  • The velocity vector of the ball.
  • The relative position vector to the opponent’s goal.
  • The relative position vector to the agent’s own goal and opponents goal.
  • Raycast measurements in the directions of North, East, West, and South, indicating the distance to the nearest boundaries of the play area.
Based on its observation, an agent selects an action. In our environment, an action is defined by five Boolean inputs: move forward, move right, move down, move left, and kick. An agent may choose any combination of these inputs, provided they are not contradictory (e.g., moving left and right simultaneously). Overall, the total number of possible actions is 18.
After an action is chosen and the simulation advances by one time step, rewards are assigned to each agent. The reward structure is designed as follows:
  • Exploration Reward: Each agent receives a small reward for moving toward the ball. This reward is computed as the dot product of the agent’s chosen direction $\vec{d}$ and the vector from the agent to the ball $\vec{v}_{a}$, scaled by a factor $\theta_{\text{exp}}$:
    $r_{\text{explore}}=\theta_{\text{exp}}\,(\vec{d}\cdot\vec{v}_{a})$ (5)
  • Team Reward: All agents on the same team receive a larger reward for moving the ball towards the opponent’s goal. This reward is calculated as the dot product of the ball’s velocity vector $\vec{v}_{b}$ and the vector from the ball to the opponent’s goal $\vec{g}$ scaled by a factor $\theta_{\text{ball}}$:
    $r_{\text{ball}}=\theta_{\text{ball}}\,(\vec{v}_{b}\cdot\vec{g})$ (6)
  • Scoring Reward: The biggest reward $r_{\text{goal}}$ is granted when a goal is scored, defined as the ball being fully inside the goal box.
  • Distance Reward: This reward is allocated based on the average distance to all teammates, to prevent players from clustering together. First, the average distance is computed, then it is capped at a maximum threshold, and finally scaled by a constant factor to prevent excessive rewards for large spacing. Specifically, if $d_{i}$ represents the distance to teammate $i$ and there are $N$ teammates, the average distance is calculated as $\bar{d}=\frac{1}{N}\sum_{i=1}^{N}d_{i}$ and the reward function is defined as:
    $f(\bar{d})=\theta_{\text{dist}}\cdot\min\left(\bar{d},\ \theta_{\text{max}}\right).$ (7)
    Here, $\theta_{\text{dist}}$ is the scaling factor and $\theta_{\text{max}}$ is the cap applied to the average distance.
It is important to note that all rewards (but the distance reward) can be negative, acting as penalties for behaviours such as moving away from the ball, moving the ball away from the opponent’s goal, or being scored on.
Games are configured to last $T$ time steps, and an episode is defined as a sequence of time steps that terminates either when a goal is scored or when the game concludes. As a result, a single game may consist of multiple episodes.
V-B Training Procedure
We employ curriculum learning [23] to progressively build up skills in this complex environment. The training is divided into the following stages:
  1. 1.
    Stage 1: Score goals against an inactive opposing team, with random ball and player spawns.
  2. 2.
    Stage 2: Introduce an opposing team with randomly selected actions, encouraging players to navigate obstacles and opponents. Continuing with random ball and player spawns.
  3. 3.
    Stage 3: Engage in league play, where a team competes against a randomly selected past version of itself while maintaining random spawn positions.
  4. 4.
    Stage 4: Continued league play with fixed starting positions for each team.
This staged learning approach allows agents to first master movement and scoring before tackling competitive play.
V-C Evaluation Metrics
Performance is evaluated using win rates, goal differentials, and ELO ratings [24]. Each team is assigned an Elo rating, which gets updated based on victory, lost or tie and the Elo rating of the opponent. When interpreting Elo ratings, it is important to note that the absolute values are less meaningful than the relative differences between them. Additionally, Goal differential is measured as the difference between the teams’ scores. In addition to performance, we assess collaboration using several key indicators:
  • Average Pairwise Distance: A higher average distance suggests that agents are not clustering excessively, which may indicate diverse task allocation.
  • Possession Frequency: This measures the number of times a ball swaps possession within the team. A high number may indicate that players are passing and, therefore, collaborating.
  • Connectivity: This metric measures the number of unobstructed graph-like connections between players, calculated as the ratio of observed connections to the maximum possible connections $\frac{N(N-1)}{2}$. To ensure only meaningful connections are counted, connections only within a range of distance are included. An optimal connectivity level indicates balanced spacing and collaborative positioning.
Using these metrics, we can comprehensively assess both the performance and collaborative behaviour of the teams.
VI Results
Our evaluation compares the performance of our TAAC algorithm against two benchmark algorithms: PPO [25] and MAAC [7]. PPO was chosen as a baseline because of its widespread use and proven robustness across diverse environments, while MAAC was chosen for being state-of-the-art in multi-agent reinforcement learning and closely aligning with our implementation. Additionally, we included a random model to provide a grounded point of reference.
We trained each algorithm for 7 days across over 100,000 games. Note that, due to its low performance and slower learning, MAAC was not advanced to stages 3 or 4 and thus did not gain experience playing against an opponent. We then assembled a league of 20 teams (5 copies of each algorithm) and randomly selected teams to compete in 5,000 games, during which we collected the following data:
Figures 4 to 6 display our performance results, illustrating the ELO ratings, goal differential, and win rates for each algorithm.
Figure 4: Elo rating for each algorithm over the evaluation period.
Figure 5: Win-rate comparison across the algorithms over the evaluation period.
Figure 6: Goal differential Matrix across the algorithms over the evaluation period. Note: The goal differential is calculated as $Score_{team_{1}}-Score_{team_{2}}$; Where a positive value indicates that team 1 is outscoring team 2 by that magnitude, while a negative value indicates the opposite.
Additionally, we present our findings on the collaborative data. Figure 7 illustrates the connectivity score for each model, Figure 8 shows the frequency of possession swaps within a team, and Figure 9 highlights the average pairwise distance among agents on the same team.
Figure 7: Connectivity comparison across the algorithms over the evaluation period.
Figure 8: Possession swaps comparison across the algorithms over the evaluation period.
Figure 9: Average Pairwise distance comparison across the algorithms over the evaluation period.
VII Analysis
Following extensive training of all algorithms, our results indicate that TAAC is the top performer, as evidenced by its superior ELO rating and win rate (see Figures 4 and 5), with PPO closely trailing. In contrast, MAAC significantly underperforms compared to expectations for a state-of-the-art MARL method, possibly due to insufficient training time or suboptimal hyperparameter settings. Additionally, Figure 6 illustrates head-to-head comparisons, highlighting that TAAC outperformed all other algorithms.
When examining collaboration-oriented metrics (Figures 79), the data suggest that TAAC exhibits more coordinated teamwork. For instance, while Figure 9 indicates that PPO agents are more spread out, Figure 7 shows that TAAC’s agents are more closely connected, which may be linked to more advantageous positioning. In addition, the possession swap metrics in Figure 8 imply that TAAC and PPO frequently swapped possession of the ball, potential indicating tactical passing. MAAC demonstrates an incredibly high possessions frequency, potentially indicating frequent passing.
Video recordings further substantiate these findings. The footage confirms that PPO frequently employs a dedicated goalie while the remaining players focus on attacking, explaining their pronounced spatial separation. Additionally, PPO exhibits clustering among its attackers and pursues the ball aggressively, which also explains its lower overall connectivity and relatively high possession frequency In contrast, TAAC consistently maintains a stable offensive formation that only transitions to a defensive position when necessary, which is evident through its connectivity score. TAAC’s dynamic strategy, characterized by frequent ball possession exchanges and timely passes, demonstrates superior tactical coordination. Video recording of MAAC indicates a similar approach to PPO, but instead having the entire team pursue the ball aggressively, this explain their low pairwise distance and connectivity. Furthermore, MAAC’s high possession frequency is also explained by this behaviour, as possession likely rapidly changes between players when the players are grouped together. Additionally, MAAC was consistently outperformed by both PPO and TAAC, as seen in the performance results.
Overall, these results suggest that our novel algorithm demonstrates enhanced collaborative behaviour and improved performance in environments where teamwork is critical. However, due to MAAC’s suboptimal performance and limited training, we plan to investigate its shortcomings further to enable a fairer comparison with state-of-the-art methods. Looking ahead, a logical next step is to evaluate TAAC in alternative environments and scenarios to better understand its limitations. Moreover, selecting an environment where collaboration is more directly tied to performance will allow us to more rigorously test our contribution.
VIII Conclusion
In this work, we introduced TAAC, a multi-agent reinforcement learning algorithm that leverages multi-headed attention in both the actor and critic networks to foster robust collaboration. Our experiments demonstrate that TAAC outperforms competing methods in win rates, goal differentials, and Elo ratings while exhibiting superior coordinated behaviour, as confirmed by game statistics and game recording analysis. Future work will involve a deeper investigation into MAAC to ensure it performs to its full potential and also to explore this new algorithm in different environments to fully test its capabilities.